OpenStudy (perii224):

Find the slope of the line passing through points P(8, -1), Q(0, -7). ¾ -4/3 -3/4 4/3

OpenStudy (anonymous):

Same method as before, you trying figuring this out on your own. @perii224

OpenStudy (anonymous):

Plug them in: (x1, y1) and (x2, y2)

OpenStudy (igreen):

\(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(y_2 = -7\) \(y_1 = -1\) \(x_2 = 0\) \(x_1 = 8\) Plug them in and solve.

OpenStudy (igreen):

Can you do that? @perii224

OpenStudy (perii224):

I Tried.. I Got -1 /:

OpenStudy (igreen):

Nope.. I got something else.

OpenStudy (igreen):

\(m =\dfrac{y_2-y_1}{x_2-x_1} \rightarrow m = \dfrac{-7 - (-1)}{0-8}\) Now can you solve?

OpenStudy (anonymous):

-6 ___ = ? and then turn the ? into a fraction and that's your answer. -8

OpenStudy (perii224):

Ok, I Got 0.75 Now

OpenStudy (igreen):

Yep, you goti t.

OpenStudy (anonymous):

What is that written as a fraction ?

OpenStudy (perii224):

So -3/4

OpenStudy (igreen):

*it

OpenStudy (anonymous):

Yes

OpenStudy (igreen):

No..not -3/4.

OpenStudy (igreen):

-3 / 4 = -0.75 not 0.75

OpenStudy (perii224):

Yup, I Got -0.75

OpenStudy (anonymous):

Oops, no. Didn't see that - sign

OpenStudy (igreen):

-0.75 is incorrect

OpenStudy (igreen):

\(m =\dfrac{y_2-y_1}{x_2-x_1} \rightarrow m = \dfrac{-7 - (-1)}{0-8} \rightarrow m = \dfrac{-6}{-8}\) Now divide that, what do you get

OpenStudy (igreen):

get?*

OpenStudy (perii224):

Oh Okay, I See Now.

OpenStudy (igreen):

Okay, what did you get as the answer?